OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (anonymous):

@johnweldon1993 helpppp

OpenStudy (johnweldon1993):

Well you know the first 2 terms are 0 and 1 ...those are given For the 3rd term...you plug 3 into that equation... f(n) = f(n - 1) + f(n - 2) f(3) = f(3 - 1) + f(3 - 2) f(3) = f(2) + f(1) f(3) = 1 + 0 f(3) = 1

OpenStudy (johnweldon1993):

and you would continue that way to find f(4) and f(5) ...but for this specific example.. knowing we have 0 , 1 , 1 for the first 3 terms...we only have 1 option

OpenStudy (anonymous):

B or C

OpenStudy (johnweldon1993):

First 3 terms are 0 , 1 , 1 B = 0 , 1 , 1, 2, 3 C = 0, 1, 2, 2, 3 So which is it? :P

OpenStudy (anonymous):

B hehe

OpenStudy (johnweldon1993):

There you go :)

OpenStudy (anonymous):

thanks hun

OpenStudy (johnweldon1993):

Anytime :)